OpenStudy (anonymous):

Raisins sell for $3.50 per pound, and granola sells for $5.90 per pound. Terri bought some raisins and some granola. The total weight was 2.1 pounds and cost $8.79. Let r represent the number of pounds of raisins Terri purchased. Which expression represents the number of pounds of granola she purchased? A. 2.1 – r B. 2.1r C. r + 2.1 D. r – 2.1

OpenStudy (anonymous):

is it c

OpenStudy (anonymous):

i dont think so, because for r+2.1, you are adding the weight of just the raisins to the total weight.

OpenStudy (anonymous):

so that would make it a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

because to find the weight of the granola, you must first subtract the weight of the raisins from the total weight.

OpenStudy (anonymous):

Raisins sell for $3.50 per pound, and granola sells for $5.90 per pound. Terri bought some raisins and some granola. The total weight was 2.1 pounds and cost $8.79. Let p represent the number of pounds of raisins. Which equation represents the situation described? A. –2.4p + 12.39 = 8.79 B. –2.4p + 8.79 = 12.39 C. 3.5p + 5.9 = 8.79 D. 3.5p + 8.79 = 5.9

OpenStudy (anonymous):

is it b

OpenStudy (anonymous):

then whatever is left would be your granola.

OpenStudy (anonymous):

plz help with this one

OpenStudy (anonymous):

hmm.

OpenStudy (anonymous):

i think it is d, i dont really have an explanation for how i got that though.

OpenStudy (anonymous):

ok i have a few more i wish i could give you 10 medals

OpenStudy (anonymous):

Apples sell for $1.90 per pound, and bananas sell for $0.75 per pound. Troy bought some apples and some bananas. Together they weighed 3.8 pounds, and cost $5.84. How many pounds of apples and how many pounds of bananas did Troy buy? A. 1.2 pounds of apples; 2.6 pound of bananas B. 1.5 pounds of apples; 2.3 pounds of bananas C. 2.6 pounds of apples; 1.2 pounds of bananas D. 1.9 pounds of apples; 1.9 pounds of bananas

OpenStudy (anonymous):

haha. no problem. i might have to go soo tho.

OpenStudy (anonymous):

no i think it's A since r is the # of the weight of raisins

OpenStudy (anonymous):

so a for the one i just put up @wolf923

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh wait i was still with the other one

OpenStudy (anonymous):

o ok lol honest mistake

OpenStudy (anonymous):

lol. :3

OpenStudy (anonymous):

so whats the one i just put up

OpenStudy (anonymous):

isa it a

OpenStudy (anonymous):

oops no a after is

OpenStudy (anonymous):

that would be c

OpenStudy (anonymous):

Soren has two jobs. During the day he works as an office clerk. In the evening he works as a cashier. His office job pays him $11 per hour. His cashier job pays him $8.75 per hour. In one week, Soren worked 46 hours. He earned a total of $470. Let c represent the number of hours that Soren works as an office clerk. Which equation can be used to find how many hours Soren worked in both jobs? A. 11c + 8.75(46 – c) = 470 B. 11c + 8.75(c – 46) = 470 C. 11c + 8.75(470 – c) = 46 D. 8.75c + 11(46 – c) = 470

OpenStudy (anonymous):

2.6x1.9=4.94 4.94+(1.2x.75)=5.84

OpenStudy (anonymous):

this is the last one, then i gotta go.

OpenStudy (anonymous):

k thanks for the help so far

OpenStudy (anonymous):

im pretty sure that one is d, but check it by inserting a number such as 2 where c is to check it.

OpenStudy (anonymous):

ok thanks for all the help to bad you could not help me on the last one

OpenStudy (anonymous):

Soren has two jobs. During the day he works as an office clerk, and in the evening he works as a cashier. His office job pays him $11 per hour. His cashier job pays him $8.75 per hour. In one week, Soren worked 46 hours. He earned a total of $470. How many hours did Soren work in each job? A. Office clerk: 16 hours; cashier: 30 hours B. Office clerk: 20 hours; cashier: 26 hours C. Office clerk: 30 hours; cashier: 16 hours D. Office clerk: 31 hours; cashier: 15 hours

OpenStudy (anonymous):

hey, im back, ill try to figure this out then i really have to go.

OpenStudy (anonymous):

your answer would be c because 30x11=330 and 330+(16x8.75)=470

OpenStudy (anonymous):

i hope you see this....